Weighted Average Cost of Capital (WACC)

Apple Inc. has 5,366,166,000 shares traded at a market value of $126 per share; its debt outstanding has an estimated market value of $90,883,140,000. The stock has a beta of 1.193. The expected return on stocks is 8.78%. The firm is rated AAA and paid 30% of its income as taxes. The risk-free rate is 2.47%, and the yield-to-maturity of Apple’s bonds is 3.27%.

A. What is Apple’s debt ratio?

B. What is Apple’s equity ratio?

C. What is Apple’s after-tax cost of debt?

D. What is Apple’s investors required return on equity?

E. What Apple’s cost of capital (after-tax WACC)?
Answer #1

(a) Number of Shares Outstanding = 5366166000 and Market Price per Share = $ 126

Market Value of Equity = 126 x 5366166000 = E = $ 676136916000

Market Value of Debt = D = $ 90883140000

Firm Value = V = D + E = $ 767020056000

Debt Ratio = D / V = 90883140000 / 767020056000 ~ 0.1185

(b) Equity Ratio = E / V = 676136916000 / 7676020056000 ~ 0.8815

(c) Yield to Maturity on Bond = 3.27 % and Tax Rate = 30 %

After-Tax Cost of Debt = kd = (1-Tax Rate) x YTM = 3.27 x (1-0.3) ~ 2.289 %

(d) Cost of Equity = Expected Return on Stock = 8.78 %

(e) WACC = Cost of Equity x Equity Ratio + After-Tax Cost of Debt x Debt Ratio = 8.78 x 0.8815 + 2.289 x 0.1185 ~ 8.011 %